OpenStudy (anonymous):

I have the answer can somone please check me.... Estimate the area under f(x)=(1/4)x^3+1 From x=0 to x=6 using 3 rectangles and left hand sums. Enter you estimates here?

OpenStudy (anonymous):

what answer did you get?

OpenStudy (anonymous):

I got 24

OpenStudy (anonymous):

@electrokid is that the right answer.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

show your steps

OpenStudy (anonymous):

ok 2 is the width of the triangles and then f(2) and F(4) and F(6) 2*2=4 2*4=8 2*6=12 4+8+12=24

OpenStudy (anonymous):

hmm... you are asked for LEFT HAND SUMS you did the right hand ones

OpenStudy (anonymous):

oh then how do you do teh left hand

OpenStudy (anonymous):

so, \[h=2\\ A=h\times\sum_{k=0}^nf(x_k)=h\times\sum_{k=0}^n\left({x_k^3\over4}+1\right) \]

OpenStudy (anonymous):

where \[x_k=x_0+k\times h\]

OpenStudy (anonymous):

ok um I am a little confused h=2 and k=0 how do I know what the rest equals

OpenStudy (anonymous):

h = 2; \[k=0,x_0=0;f(x_0)=1\] \[k=1,x_1=0+2=2,f(x_1)={2^3\over4}+1=?\] \[k=2,x_2=0+4=4,f(x_2)={4^3\over4}+1=?\] \[A=h\times[f(x_0)+f(x_1)+f(x_2)]\]

OpenStudy (anonymous):

42 would be the answer then

OpenStudy (anonymous):

@electrokid is that the answer 42

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then should the estimate be over or under the true estimate? is the second part of the question and I can't figure it out I think it should be under

OpenStudy (anonymous):

do you know which it should be

OpenStudy (anonymous):

is the function continuously increasing or decreasing? you can check with the first derivative.

OpenStudy (anonymous):

if the function is continuously increasing, the left hand approximation is an "under approximation" and vice versa

OpenStudy (anonymous):

ok yes it is increasing so then it is under correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@farmergirl411 you are supposed to give a medal when you are helped